# EpiDiff: Enhancing Multi-View Synthesis via Localized Epipolar-Constrained Diffusion

### v2

@@ 1 Introduction
Recently, with the introduction of diffusion models [17, 41], the field of 3D generation has experienced rapid and vigorous development. A series of research [30, 34, 68, 11, 28, 1, 5, 7, 12, 15, 21, 33, 35, 58, 20, 67, 2] attempts to train 3D generative diffusion models from scratch on 3D assets. However, due to the limited size of publicly available 3D datasets, these methods demonstrate poor generalizability. Therefore, another group of works [18, 36, 55, 6, 60, 31, 37, 38, 50, yu2023csd, sun2023dreamcraft3d] propose to distill prior knowledge of 2D diffusion models to create 3D models from text or images via Score Distillation Sampling (SDS) [36]. Although these methods have achieved impressive results, they require expensive per-scene optimization that usually takes tens of minutes or even hours.
@@
In this paper, we present EpiDiff, a novel localized interactive multiview diffusion model. The key idea is to utilize the feature maps of neighboring views that suffer fewer occlusions for the generation of target view images, which improves the generalization capability and efficiency. To fuse the neighboring feature maps, during the generation of $N$ target views, we insert a lightweight epipolar attention block into the UNet, which leverages epipolar constraints to enable interaction among feature maps of $K$ neighboring views within a smaller range. The 3D modeling module preserves the original feature distribution of the diffusion model, exhibiting compatibility with various base diffusion models. Compared to global modeling methods, our localized modeling approach not only speeds up the generation process, but also facilitating the generation of more diversely distributed views, contributing to improved reconstruction from generated multiviews.
@@ 2 Related work
3D generation guided by 2D prior  Given the robust generative capabilities of 2D diffusion models [41, 39, 42], some research efforts directly leverage pretrained image diffusion models. With DreamFusion [36] and SJC [55] proposing to distill a pretrained 2D text-to-image generation model for generating 3D shapes, a series of subsequent studies [23, 6, 60, 31, 37, 38, 50, yu2023csd, sun2023dreamcraft3d] adopt such a pre-scene optimization pipeline. These methods optimize a 3D representation (e.g., NeRF, mesh, or SDF) conditioned on text or images through Score Distillation Sampling (SDS) [36]. However, the per-scene optimization approach has issues with efficiency. Generating a single scene may take tens of minutes or even hours.
Using 2D diffusion for multiview synthesis  2D diffusion models have introduced significant progress in generating multiview images from a single view. While some approaches [3, 9, 13, 51, 54, 61, 62, 66] integrate 3D representations for multiview synthesis, they lack design considerations for 3D reconstruction. Zero123 [25] introduces relative viewpoint condition to 2D diffusion models, enabling novel view synthesis (NVS) from a single image. However, due to its limited consistency, it utilizes SDS loss [36] to achieve 3D reconstruction. One-2-3-45 [24] leverages a generalizable neural reconstruction method [29] for directly producing 3D shapes from multiview images generated by Zero123 [25], yet the resultant quality is suboptimal. More recently, SyncDreamer [27] has introduced a 3D global feature volume during multiview generation, to ensure consistency. However, the heavy global 3D representation modeling not only slows down the generation but also tends to produce low-quality images. In contrast, our localized modeling method facilitates the rapid generation of multiview-consistent and high-quality images.
Generalizable radiance field  Since NeRF [32] achieved advancements in fitting single scene for novel view synthesis, numerous studies have explored generalizable neural radiance field representation for various scenes. These methods take a few input view images and generate a 3D implicit field in a single feed-forward pass, eliminating the need for per-scene optimization. Among the methods, some explicitly construct the 3D feature or cost volume [4, 19, 70, 29] and utilize the voxel feature for decoding density and color, inspiring SyncDreamer[27] to model 3D global representation, while others [57, 40, 16, 64, 63, 26, 22, 52, 43, 44, 49] directly aggregate 2D features using MLPs or transformers, guided by spatial projection relationships. These methods, either explicitly or implicitly modeling generalized scenes, provide valuable insights for consistency modeling in the feed-forward multiview generation process.
@@ 3 Method
To enhance the consistency, quality and efficiency of the multiview generation, we introduce a localized interactive multiview diffusion scheme based on the novel view synthesis base model. We design a lightweight attention block in the original feature space of the diffusion model, with the aim of establishing nearby view associations to model 3D representation. As illustrated in Fig. 2, the proposed approach involves an interaction process among the intermediate features of multiview images (refer to Sec. 3.1), thereby extending single-view synthesis to multiview synthesis. Specifically, we design a 3D modeling module, named Epipolar-constrained Attention Block (ECA Block) based on the attention mechanism and insert it into the base diffusion model (refer to Sec. 3.2). The ECA Block facilitates information exchange among latent features of neighboring views, leveraging epipolar constraints in stereo vision to enhance consistency and quality in multiview images. The ECA Block, with its lightweight and highly adaptable characteristics, is suitable for various base models.
@@ 4.1 Experiment settings
We evaluate EpiDiff using the Google Scanned Object dataset [10], which consists of over one thousand scanned models, distinct from the training distribution. To assess multiview synthesis, we employ two rendering settings. Firstly, aligning with SyncDreamer’s training setting [27], we fix elevation at $30^{\circ}$ and sample 16 cameras with azimuths evenly spread from $0^{\circ}$ to $360^{\circ}$. Secondly, we render 16 views with azimuths also uniformly distributed in $[0^{\circ},360^{\circ}]$, but with varying elevations from $-10^{\circ}$ to $40^{\circ}$. This tests the model’s ability to generate multiview images across a wider range of viewpoints.
@@ 4.4 Ablation study
To assess the effectiveness of our design for ECA Block (Sec. 3.2), we conducted ablation experiments on multiview synthesis at elevation $30^{\circ}$ setting (Tab. 4). We investigated the impact of the number of neighboring views used for aggregating features in Near-Views Cross Attention module, setting $F$ to $\{4,8,16\}$, where $K=16$ indicates the use of all perspective features aggregated to the target view. Experimental results show that increasing $F$ consumes more time and reduces the quality of the generated multiview structure and appearance. Additionally, we examined the positional encoding (PE) strategies and found that removing either the light field PE or the ray-relative PE leads to a decrease in the consistency and quality of the generated images.

### v3

@@ 1 Introduction
Recently, with the introduction of diffusion models [21, 45], the field of 3D generation has experienced rapid and vigorous development. A series of research [34, 38, 70, 14, 32, 1, 6, 9, 15, 19, 25, 37, 39, 59, 24, 67, 3] attempts to train 3D generative diffusion models from scratch on 3D assets. However, due to the limited size of publicly available 3D datasets, these methods demonstrate poor generalizability. Therefore, another group of works [22, 40, 56, 7, 61, 35, 41, 42, 52, 69, 51] propose to distill prior knowledge of 2D diffusion models to create 3D models from text or images via Score Distillation Sampling (SDS) [40]. Although these methods have achieved impressive results, they require expensive per-scene optimization that usually takes tens of minutes or even hours.
@@
In this paper, we present EpiDiff, a novel localized interactive multiview diffusion model. The key idea is to utilize the feature maps of neighboring views that suffer fewer occlusions for the generation of target view images, which improves the generalization capability and efficiency. To fuse the neighboring feature maps, during the generation of $N$ target views, we insert a lightweight epipolar attention block into the UNet, which leverages epipolar constraints to enable interaction among feature maps of $F$ neighboring views within a smaller range. The 3D modeling module preserves the original feature distribution of the diffusion model, exhibiting compatibility with various base diffusion models. Compared to global modeling methods, our localized modeling approach not only speeds up the generation process, but also facilitating the generation of more diversely distributed views, contributing to improved reconstruction from generated multiviews.
@@ 2 Related work
3D generation guided by 2D prior  Given the robust generative capabilities of 2D diffusion models [45, 43, 46], some research efforts directly leverage pretrained image diffusion models. With DreamFusion [40] and SJC [56] proposing to distill a pretrained 2D text-to-image generation model for generating 3D shapes, a series of subsequent studies [27, 7, 61, 35, 41, 42, 52, 69, 51, 18, 68, 2] adopt such a pre-scene optimization pipeline. These methods optimize a 3D representation (e.g., NeRF, mesh, or SDF) conditioned on text or images through Score Distillation Sampling (SDS) [40]. However, the per-scene optimization approach has issues with efficiency. Generating a single scene may take tens of minutes or even hours.
Using 2D diffusion for multiview synthesis  2D diffusion models have introduced significant progress in generating multiview images from a single view. While some approaches [4, 11, 16, 53, 55, 62, 63, 66, 13] integrate 3D representations or utilize shared attention for multiview or texture synthesis, they lack design for 3D reconstruction. Zero123 [29] introduces relative view condition to diffusion models, enabling novel view synthesis (NVS) from a single image. However, due to its limited consistency, it utilizes SDS loss [40] to achieve 3D reconstruction. One-2-3-45 [28] leverages a generalizable neural reconstruction method [33] for producing 3D shapes from multiview images generated by Zero123 [29], yet the resultant quality is suboptimal. More recently, SyncDreamer [31] has introduced a 3D global feature volume during multiview generation to ensure consistency. However, the heavy global 3D representation modeling not only slows down the generation but also tends to produce low-quality images. In contrast, our localized modeling method facilitates the generation of multiview-consistent and high-quality images.
Generalizable radiance field  Since NeRF [36] achieved advancements in fitting single scene for novel view synthesis, numerous studies have explored generalizable neural radiance field representation for various scenes. These methods take a few input view images and generate a 3D implicit field in a single feed-forward pass, eliminating the need for per-scene optimization. Among the methods, some explicitly construct the 3D feature or cost volume [5, 23, 72, 33] and utilize the voxel feature for decoding density and color, inspiring SyncDreamer[31] to model 3D global representation, while others [58, 44, 20, 65, 64, 30, 26, 54, 47, 48, 50, 8] directly aggregate 2D features using MLPs or transformers, guided by spatial projection relationships. These methods, either explicitly or implicitly modeling generalized scenes, provide valuable insights for consistency modeling in the feed-forward multiview generation process.
@@ 3 Method
To enhance the consistency, quality and efficiency of the multiview generation, we introduce a localized interactive multiview diffusion scheme based on the novel view synthesis base model. We design a lightweight attention block in the original feature space of the diffusion model, with the aim of establishing nearby view associations to model 3D representation. As illustrated in Fig. 2, the proposed approach involves an interaction process among the intermediate features of multiview images (refer to Sec. 3.1), thereby extending single-view synthesis to multiview synthesis. Specifically, we design a 3D modeling module, named Epipolar-constrained Attention Block (ECA Block) based on the attention mechanism and insert it into the base diffusion model (refer to Sec. 3.2). The ECA Block facilitates information exchange among latent features of neighboring views, leveraging epipolar constraints in stereo vision to enhance consistency and quality in multiview images. The ECA Block, with its lightweight and adaptable characteristics, is suitable for various base models.
@@ 4.1 Experiment settings
We evaluate EpiDiff using the Google Scanned Object dataset [12], which consists of over one thousand scanned models. To assess multiview synthesis, we employ two rendering settings. Firstly, aligning with SyncDreamer’s training setting [31], we fix elevation at $30^{\circ}$ and sample 16 cameras with azimuths evenly spread from $0^{\circ}$ to $360^{\circ}$. Secondly, we render 16 views with azimuths also uniformly distributed in $[0^{\circ},360^{\circ}]$, but with varying elevations from $-10^{\circ}$ to $40^{\circ}$. This tests the model’s ability to generate multiview images across a wider range of viewpoints.
@@ 4.4 Ablation study
To assess the effectiveness of our design for ECA Block (Sec. 3.2), we conducted ablation experiments on multiview synthesis at elevation $30^{\circ}$ setting (Tab. 4). We investigated the impact of the number of neighboring views used for aggregating features in Near-Views Cross Attention module, setting $F$ to $\{4,8,16\}$, where $F=16$ indicates the use of all perspective features aggregated to the target view. Results show that increasing $F$ consumes more time and reduces the quality of the generated multiview structure and appearance. Additionally, we examined the positional encoding (PE) strategies and found that removing either the light field PE or the ray-relative PE leads to a decrease in the consistency and quality of the generated images.
@@ 5 Limitations and Conclusion
Acknowledgment
This work was supported by National Natural Science Foundation of China (62132001).